OpenStudy (anonymous):

I have a Pre-Calc Test Monday, and our teacher told us one of the questions is sinx^2+cosx^2+1=0 or sinx^2+cosx+1=0. SOLVE FOR GENERAL SOLUTIONS, please! I am not sure which one is the correct problem, so could someone help me solve both? Thanks so much!

OpenStudy (phi):

is that \[ \left(\sin(x)\right)^2 = \sin^2(x)\] or \[ \sin\left(x^2\right) \]

OpenStudy (anonymous):

It is the first set (sin(x))2=sin2(x)

OpenStudy (phi):

in that case, you should remember this useful fact \[ \sin^2(x) + \cos^2(x) = 1 \]

OpenStudy (phi):

if you use that identity in the first equation you gave, you get 2=0 which is not true.

OpenStudy (phi):

you can use it in the second equation you have, to get a quadratic in cos(x)

OpenStudy (anonymous):

Wouldn't the 1 be -1 in the original equation if I was using the identity sin2(x)+cos2(x)=1 because it would have been moved over to the other side?

OpenStudy (phi):

\[ \sin^2x+\cos^2x+1=0\] use the identity to replace the first 2 terms with 1. you get 1+1 = 0

OpenStudy (anonymous):

Oh Okay! I see now!

OpenStudy (phi):

for the 2nd one \[ \sin^2x+\cos x+1=0 \\ (1 -\cos^2 x ) + \cos x+1=0 \] that simplifies to \[ -\cos^2 x + \cos x +2 = 0 \\ \cos^2 x - \cos x - 2 = 0 \] if you let y = cos x, this is \[ y^2 -y -2 = 0 \] factor to get \[ (y-2)(y+1) = 0 \] or (y-2) =0 --> cos x = 2 Not possible (cos goes -1 to +1) y+1=0 --> cos x = -1 now find all angles x where cos x is -1

OpenStudy (anonymous):

Thank you so much!!!!!!